OpenStudy (anonymous):

How do I find the value of i^(1/4)?

OpenStudy (anonymous):

I know you have to turn it to polar form

OpenStudy (diamondboy):

x^1/4 is the same as d fourth root of X

OpenStudy (anonymous):

i dont get what your saying @diamondboy

OpenStudy (freckles):

\[i=\cos(\frac{\pi}{2}+2n \pi)+i \sin(\frac{\pi}{2}+2n \pi) \text{ where } n \text{ is an integer } \]

OpenStudy (freckles):

now raise both sides to 1/4 power

OpenStudy (anonymous):

what do i put for n?

OpenStudy (anonymous):

how did you get the theta?

OpenStudy (freckles):

n=0,1,2,3 you will give the 4 answers you seek I just recalled that cos(pi/2)=0 and sin(pi/2)=1

OpenStudy (freckles):

i=a+bi when a=0 and b=1 right?

OpenStudy (freckles):

and cos is 0 when theta=pi/2 and sin is 1 when theta=pi/2

OpenStudy (anonymous):

we never had to do that, so i dont want to do something different

OpenStudy (diamondboy):

are u talking about complex numbers here @_alex_urena_

OpenStudy (anonymous):

let z=i^1/4 \[z^4=i=r(\cos \frac{ \pi }{ 2 }+i \sin \frac{ \pi }{ 2 })=r e ^{i \frac{ \pi }{ 2 }}=re ^{i \left( \frac{ \pi }{ 2 }+2n \pi \right)}\] \[z^4=r e ^{i\left( 4n+1 \right)\frac{ \pi }{ 2 }}\] here r=1 \[z=e ^{i \left( 4n+1 \right)\frac{ \pi }{ 8 }},n=0,1,2,3\] ?

OpenStudy (anonymous):

yass, powers and roots of complex numbers

OpenStudy (anonymous):

we never did something like that @surjithayer

OpenStudy (diamondboy):

ok so what if u did this